OpenStudy (anonymous):

The length of a field is 3 times its width. If the perimeter (distance around the field) of the field is 312 feet, what is the width of the field?

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (skullpatrol):

Let w = the width of the field in feet

OpenStudy (anonymous):

umm 3w=312?

OpenStudy (skullpatrol):

Nope. Then The length of a field is 3 times its width. means 3•w = length of the field

OpenStudy (skullpatrol):

Does that^ make sense?

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

Recall the formula for the perimeter of a rectangle is P = 2l + 2w

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what would be the width ?

OpenStudy (skullpatrol):

The length of a field is 3 times its width. If the perimeter (distance around the field) of the field is 312 feet, what is the width of the field? P = 2l + 2w 312 = 2( 3•w ) + 2w

OpenStudy (anonymous):

where'd you get the two

OpenStudy (skullpatrol):

Which 2?

OpenStudy (anonymous):

2(3*w)

OpenStudy (anonymous):

wy do you multiply it by 2

OpenStudy (skullpatrol):

It is in the formula right above it.

OpenStudy (skullpatrol):

P = 312 and l = 3•w

OpenStudy (anonymous):

oooh stupid me haha okay now I get it

OpenStudy (skullpatrol):

Now, solve the equation for w.

OpenStudy (anonymous):

I got 76.5

OpenStudy (anonymous):

2*3*w = 6*2w?

OpenStudy (anonymous):

ooh no 6w+2w= 8w so 312/8= 39

OpenStudy (anonymous):

thanks :)

OpenStudy (skullpatrol):

Great job :D

OpenStudy (skullpatrol):

Don't forget the units of feet on your answer.